Camphor (C10H16O) has a molal freezing point depression constant of 5.95°C/m. If you dissolve 10.0 g of dimethyl ether (C2H6O) into 3.00 kg of camphor, what will be the change in the freezing point of camphor?

1
Expert's answer
2022-04-15T05:02:03-0400

Solution:

solute = dimethyl ether (C2H6O)

solvent = camphor (C10H16O)


The equation describing the change in freezing point from pure solvent to solution is:

Δt = i × Kf × m

where:

Δt = the change in freezing point (°C)

i = the van't Hoff factor

Kf = the molal freezing point depression constant (°C m−1)

m = the molality of the solute (m)


The molar mass of dimethyl ether (C2H6O) is 46.07 g mol−1

Therefore,

Moles of C2H6O = (10.0 g C2H6O) × (1 mol C2H6O / 46.07 g C2H6O) = 0.217 mol C2H6O

Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of C2H6O solution = (0.217 mol) / (3.00 kg) = 0.0723 mol/kg = 0.0723 m

 

i = 1 (for nonelectrolyte - dimethyl ether)

Kf = 5.95°C m−1 (for camphor)


Thus,

Δt = (1) × (5.95°C m−1) × (0.0723 m) = 0.43°C

Δt = 0.43°C

 

Answer: The change in the freezing point of camphor will be 0.43°C